OpenStudy (anonymous):

one canned juice drink is 20% orange juice; another is 5% orange juice. how many liters of each should be mixed together in order to get 15 L that is 8%

OpenStudy (chosenmatt):

take away 16 I think @WolfMaster @CausticSyndicalist @Kit_Kat_Nat_<3 right guys?

OpenStudy (dumbcow):

x + y = 15 20x + 5y = 15(8) solve the system